# SemTalk: Holistic Co-speech Motion Generation with Frame-level Semantic Emphasis

### v1

@@ Abstract
A good co-speech motion generation cannot be achieved without a careful integration of common rhythmic motion and rare yet essential semantic motion. In this work, we propose SemTalk for holistic co-speech motion generation with frame-level semantic emphasis. Our key insight is to separately learn general motions and sparse motions, and then adaptively fuse them. In particular, rhythmic consistency learning is explored to establish rhythm-related base motion, ensuring a coherent foundation that synchronizes gestures with the speech rhythm. Subsequently, semantic emphasis learning is designed to generate semantic-aware sparse motion, focusing on frame-level semantic cues. Finally, to integrate sparse motion into the base motion and generate semantic-emphasized co-speech gestures, we further leverage a learned semantic score for adaptive synthesis. Qualitative and quantitative comparisons on two public datasets demonstrate that our method outperforms the state-of-the-art, delivering high-quality co-speech motion with enhanced semantic richness over a stable base motion.
@@ 1 Introduction
Nonverbal communication, including body language, hand gestures, and facial expressions, is integral to human interactions. It enriches conversations with contextual cues and enhances understanding among participants [13, 19, 23, 6]. This aspect is particularly significant in holistic co-speech motion generation, where the challenge lies in synthesizing gestures that align with speech rhythm while also capturing the infrequent yet critical semantic gestures that convey specific ideas [37, 24].
@@
Most existing methods [47, 29, 16] rely heavily on rhythm-related audio features as conditions for gesture generation. While these rhythm-based features successfully align gestures with the timing of speech, they often overshadow the sparse yet expressive semantic motion (see Fig. 1). As a result, the generated motions may lack the contextual depth necessary and nuanced expressiveness for natural interaction. Some methods try to address this by incorporating semantic information like emotion, style, and content[9, 11, 22, 31]. However, rhythm features tend to dominate, making it difficult to capture sparse, semantically relevant gestures at the frame level. These rare but impactful gestures are often diluted or overlooked, highlighting the challenge of balancing rhythmic alignment with semantic expressiveness in co-speech motion generation.
@@
In real-world human conversations, we have an observation that while most speech-related gestures are indeed rhythm-related, only a limited number of frames involve semantically emphasized gestures. This insight suggests that co-speech gestures can be decomposed into two distinct components: ($i$) General rhythm-related motions. These provide a continuous, coherent base motion aligned with the speech rhythm, reflecting the natural timing of speaking. ($ii$) Sparse semantic-aware motions: These occur infrequently but are essential for conveying specific meanings or emphasizing key points within the conversation.
@@
Inspired by this observation, we propose a new framework SemTalk. SemTalk models the base motion and the sparse motion separately and then fuse them adaptively to generate high-fidelity co-speech motion. Specifically, we first focus on capturing the foundational patterns of speech by introducing rhythmic consistency learning. This method aligns the generated motions with rhythm by a fine-grained refinement and local-global consistency learning. For fine-grained refinement, we incorporate a coarse2fine cross-attention module, which progressively refines the base motion cues in a coarse-to-fine manner, starting from the face and moving through the hands, upper body, and lower body. This approach ensures consistent rhythmic transmission across all body parts, enhancing coherence and rhythmic base motion. For local-global consistency learning, locally, we use a frame-level consistency loss that encourages each frame to synchronize with the corresponding speech features, ensuring precise temporal alignment. Globally, we employ a sequence-level consistency loss that maintains the overall rhythmic flow across the entire motion sequence.
@@
Furthermore, we introduce semantic emphasis learning, which focuses on generating sparse yet meaningful semantic motion. This approach utilizes frame-level semantic cues from textual information, high-level speech features, and emotion labels to identify frames that require emphasis. A gating mechanism is then employed to adaptively activate semantic motions at key frames, allowing the model to produce gestures that not only align with the rhythm but also enhance the motion with deeper semantic meaning and contextual relevance. Finally, the learned base motion and sparse motion are integrated through a dynamically learned semantic score, which selectively amplifies expressiveness by incorporating semantic gestures into the rhythm-aligned base motion. This adaptive approach enables the model to capture the full spectrum of human expressiveness, generating motions that are semantically aligned with the speech.
@@
  • We propose SemTalk, a novel framework for holistic co-speech motion generation that separately models general rhythm-related base motion and semantic-aware sparse motion, adaptively integrating them based on a learned semantic score.
  • We introduce rhythmic consistency learning to incorporate latent face and latent hands features with rhythm-related priors, ensuring coherent and rhythm-related base motion. We then propose semantic emphasis learning to synthesize semantic gestures at certain frames, enhancing expressive semantic-aware sparse motion.
@@ 2 Related Work
Co-speech Gesture Generation. Co-speech gesture generation focuses on creating gestures that naturally align with speech to enhance communication. Early rule-based methods [20, 21, 7, 18, 40] lacked the ability to capture the variability of human motion. Deterministic models [48, 45, 5, 7, 35] attempted direct mappings from speech to gestures. Probabilistic models like GANs [16, 39, 1] and diffusion models [52, 2, 46, 9] introduced variability. Some approaches incorporated semantic information. HA2G [31] and SEEG [27] used hierarchical networks and alignment techniques. LivelySpeaker [51] focuses on generating co-speech gestures by conditioning on both rhythmic features and semantic information. However, it lacked adaptability for integrating semantic gestures with the speech rhythm. DisCo [28] also addresses gesture diversity by disentangling content and rhythm, and balancing data distributions for more diverse motions, but it does not explicitly model sparse semantic gestures or dynamically integrate them with the rhythmic base motion. In contrast, SemTalk introduces a novel framework that separately models rhythm-related base motions and generates sparse semantic gestures. To ensure two components are seamlessly integrated, we dynamically combine base motion and sparse motion based on a learned semantic score.
Holistic Co-speech Motion Generation. Generating synchronized, expressive full-body motion in response to speech remains challenging, especially in coordinating the face, hands, and torso [33, 47, 30, 8, 36]. Previous methods gave way to generative models that enhance synchronization across regions. Habibie et al. [16] introduced a CNN-based approach for face and body generation, though it often led to uncoordinated gestures. TalkSHOW [47] improved synchronization with VQ-VAE [41] cross-conditioning but handled facial expressions separately, causing fragmented outputs. DiffSHEG [8] and EMAGE [30] partially address this by using separate encoders for expressions and gestures, though unidirectional flow restricts expressive coherence, and ProbTalk [32] leverages PQ-VAE [42] for better body-facial synchronization. These methods, however, primarily rely on audio-related rhythmic cues, risking the loss of nuanced semantic gestures. Inspired by TM2D [14], which decomposes dance motion into music-related and text-aware components, we decompose co-speech motion into rhythm-related and sparse semantic motions, learning each independently. This separation enables us to better capture both the rhythmic flow and the expressive nuances of the motion.
@@ 3.1 Preliminary on RVQ-VAE
Following [50, 4, 15], our approach uses a residual vector-quantized autoencoder (RVQ-VAE) to precisely capture complex body movements in a latent space. Unlike standard VQ-VAE [41], which has a single quantized latent space, RVQ-VAE decomposes motion into multiple quantization levels, capturing both coarse and fine details. To retain unique motion characteristics across body regions, we segment the body into four parts—face, upper body, hands, and lower body—each with a dedicated RVQ-VAE, following [3, 30]. This segmentation preserves each part’s dynamics and prevents feature entanglement. Each part’s motion sequence, denoted $M=m_{1:N}\in\mathbb{R}^{N\times D}$, is mapped to a latent sequence $z_{1:n}\in\mathbb{R}^{n\times d}$ via a 1D convolutional encoder with downsampling factor $n/N$. This latent sequence undergoes a series of quantization steps $Q_{1},Q_{2},\ldots,Q_{V}$ to progressively capture fine details. Each quantization layer assigns the latent vector to the nearest codebook entry, recording residuals at each step. This cumulative encoding yields a refined latent representation $\hat{z}$, and the decoding mirrors the encoding, with distinct decoders for each region.
@@ 3.2 Overview
As shown in Figure 2, our SemTalk pipeline includes two main components: the Base Motion Blocks $f_{r}(\cdot)$ and the Sparse Motion Blocks $f_{s}(\cdot)$. Given rhythmic features $\gamma_{b}$, $\gamma_{h}$, a seed pose $\tilde{m}$, and a speaker ID $id$, the Base Motion Blocks generate rhythm-aligned codes $q^{r}$, forming the rhythmic foundation of the base motion:
$f_{r}:(\gamma_{b},\gamma_{h},\tilde{m},id;\theta_{f_{r}})\rightarrow q^{r},$ (1)
@@
where $\theta_{f_{s}}$ represents the Sparse Motion Block parameters. Finally, the semantic emphasis mechanism $\mathcal{E}$ combines $q^{r}$ and $q^{s}$, guided by $\psi$, to form the final motion codes $q^{m}$:
$q^{m}=\mathcal{E}(q^{r},q^{s};\psi).$ (3)
@@
Figure 3: Architecture of SemTalk. SemTalk generates holistic co-speech motion in three stages. (a) Base Motion Generation uses rhythmic consistency learning to produce rhythm-aligned codes $q^{r}$, conditioned on rhythmic features $\gamma_{b}$, $\gamma_{h}$, seed pose $\tilde{m}$, and $id$. (b) Sparse Motion Generation employs semantic emphasis learning to generate semantic codes $q^{s}$, activated by semantic score $\psi$. (c) Adaptively Fusion automatically combines $q^{r}$ and $q^{s}$ based on $\psi$ to produce mixed codes $q^{m}$ at frame level for rhythmically aligned and contextually rich motions.
@@ 3.3 Generating Rhythm-related Base Motion
The Base Motion Generation (Fig. 3 a) in SemTalk establishes a rhythmically aligned foundation for co-speech gesture generation by leveraging both rhythmic and speaker-specific features, enhancing the naturalness and personalization of generated gestures.
Rhythmic Speech Encoding. To synchronize motion with speech, SemTalk incorporates rhythmic features: beats $\gamma_{b}$ and HuBERT features $\gamma_{h}$. $\gamma_{b}$, derived from amplitude, short-time energy [10], and onset detection, mark key rhythmic points for aligning gestures with speech. Meanwhile, $\gamma_{h}$, extracted by the HuBERT encoder [17], captures high-level audio traits such as phonetic structure and speaker nuances, adding details to the gestures. In addition to rhythmic features $\gamma$, SemTalk uses a seed pose $\tilde{m}$ and speaker identity $id$ to generate a personalized, rhythm-aligned latent pose $p$. Then MLP-based Face Enhancement and Body Part-Aware modules utilize $\gamma$, $p$ and $id$ to obtain latent face $f_{e}$, hands $f_{h}$, upper body $f_{u}$ and lower body $f_{l}$.
Rhythmic Consistency Learning. To enforce a rhythmically coherent base motion, we introduce rhythmic consistency learning, which includes a fine-grained refinement and local-global consistency learning designed to align generated gestures with the speech. To refine base motion, we first proposed a transformer-based Coarse2Fine Cross-Attn Module utilize $\gamma$, $p$ and $id$ to obtain latent base motion $f_{b}$. The refinement begins with $\gamma$ for $f_{e}$, which guides the rhythmic representation for $f_{h}$, followed by conditioning $f_{u}$ and finally influencing $f_{l}$. This progression from face to hands, upper body, and lower body enables the model to create synchronized motions that align naturally with spoken content. An MLP layer then generates codes that synchronize different body regions with the rhythm, producing distinct codes for the face $q^{r}_{f}$, hands $q^{r}_{h}$, upper body $q^{r}_{u}$, and lower body $q^{r}_{l}$, ensuring temporal coherence across all parts.
Inspired by CoG’s use of InfoNCE loss [44] to synchronize facial expressions with audio cues, our approach adopts a similar philosophy of aligning motion and speech. However, we differ fundamentally by incorporating separate local and global rhythmic consistency losses, which are applied to both latent face $f_{e}$ and latent hands $f_{h}$, ensuring a more cohesive and synchronized representation across the entire motion sequence. InfoNCE loss maximizes mutual information and ensures temporal alignment by contrasting frame-level or sentence-level features across modalities, synchronizing rhythm. By leveraging HuBERT features $\gamma_{h}$ instead of basic beat features $\gamma_{b}$, which only capture rhythmic pauses, we incorporate rich, high-level audio representations that enhance the model’s ability to capture rhythm-related motion patterns and maintain temporal coherence.
This rhythmic consistency loss ensures that the motions are not only synchronized at the frame level but also maintain a consistent rhythmic flow across the entire sequence. It can be defined as:
@@
The local frame-level consistency loss $\mathcal{L}_{\text{Rhy}}^{(L)}$ aligns the motion features of each frame with the corresponding rhythmic cues. This local loss encourages each frame’s gestures and facial movements to align with subtle rhythmic variations, embedding rhythm within each temporal segment.
To ensure rhythmic coherence at a broader level, we incorporate a global sample-level consistency loss $\mathcal{L}_{\text{Rhy}}^{(G)}$. Unlike local loss, $\mathcal{L}_{\text{Rhy}}^{(G)}$ reinforces rhythm consistency throughout the sequence, ensuring that the generated motion maintains a smooth, rhythm-aligned motion flow throughout its duration. By jointly minimizing $\mathcal{L}_{\text{Rhy}}^{(L)}$ and $\mathcal{L}_{\text{Rhy}}^{(G)}$, rhythmic consistency learning enables SemTalk to produce base motions that are rhythmically aligned and temporally cohesive, forming a solid foundation upon which further semantic gesture refinement can be applied.

### v3

@@ Abstract
A good co-speech motion generation cannot be achieved without a careful integration of common rhythmic motion and rare yet essential semantic motion. In this work, we propose SemTalk for holistic co-speech motion generation with frame-level semantic emphasis. Our key insight is to separately learn base motions and sparse motions, and then adaptively fuse them. In particular, coarse2fine cross-attention module and rhythmic consistency learning are explored to establish rhythm-related base motion, ensuring a coherent foundation that synchronizes gestures with the speech rhythm. Subsequently, semantic emphasis learning is designed to generate semantic-aware sparse motion, focusing on frame-level semantic cues. Finally, to integrate sparse motion into the base motion and generate semantic-emphasized co-speech gestures, we further leverage a learned semantic score for adaptive synthesis. Qualitative and quantitative comparisons on two public datasets demonstrate that our method outperforms the state-of-the-art, delivering high-quality co-speech motion with enhanced semantic richness over a stable base motion.
@@ 1 Introduction
Nonverbal communication, including body language, hand gestures, and facial expressions, is integral to human interactions. It enriches conversations with contextual cues and enhances understanding among participants [14, 20, 24, 6]. This aspect is particularly significant in holistic co-speech motion generation, where the challenge lies in synthesizing gestures that align with speech rhythm while also capturing the infrequent yet critical semantic gestures [38, 25].
@@
Most existing methods [48, 30, 17] rely heavily on rhythm-related audio features as conditions for gesture generation. While these rhythm-based features successfully align gestures with the timing of speech, they often overshadow the sparse yet expressive semantic motion (see Fig. 1). As a result, the generated motions may lack the contextual depth necessary and nuanced expressiveness for natural interaction. Some methods try to address this by incorporating semantic information like emotion, style, and content[10, 12, 23, 32]. However, the rhythm features tend to dominate, making the models difficult to capture sparse, semantically relevant gestures at the frame level. These rare but impactful gestures are often diluted or overlooked, highlighting the challenge of balancing rhythmic alignment with semantic expressiveness in co-speech motion generation.
@@
In real-world human conversations, we have an observation that while most speech-related gestures are indeed rhythm-related, only a limited number of frames involve semantically emphasized gestures. This insight suggests that co-speech motions can be decomposed into two distinct components: ($i$) Rhythm-related base motion. These provide a continuous, coherent base motion aligned with the speech rhythm, reflecting the natural timing of speaking. ($ii$) Semantic-aware sparse motion: These occur infrequently but are essential for conveying specific meanings or emphasizing key points within the conversation.
@@
Inspired by this observation, we propose a new framework SemTalk. SemTalk models the base motion and the sparse motion separately and then fuses them adaptively to generate high-fidelity co-speech motion. Specifically, we first focus on generating rhythm-related base motion by introducing coarse2fine cross-attention module and rhythmic consistency learning. We design a hierarchical coarse2fine cross-attention module, which progressively refines the base motion cues in a coarse-to-fine manner, starting from the face and moving through the hands, upper body, and lower body. This approach ensures consistent rhythmic transmission across all body parts, enhancing coherence base motion. Moreover, we propose a local-global rhythmic consistency learning approach, which enforces alignment at both the frame and sequence levels. Locally, a frame-level consistency loss ensures that each frame is precisely synchronized with its corresponding speech features, guaranteeing accurate temporal alignment. Globally, a sequence-level consistency loss sustains a coherent rhythmic flow across the entire motion sequence, preserving consistency throughout the generated gestures.
@@
Furthermore, we introduce semantic emphasis learning approach, which focuses on generating semantic-aware sparse motion. This approach utilizes frame-level semantic cues from textual information, high-level speech features, and emotion to identify frames that require emphasis through a learned semantic score produced by a gating strategy, i.e., sem-gate. The sem-gate is designed to dynamically activate semantic motions at key frames through two weighting methods applied on the motion condition and the loss, respectively, and semantic label guidance, allowing the model to produce motion that enhances the motion with deeper semantic meaning and contextual relevance.
Finally, the base motion and sparse motion are integrated through semantic score-based motion fusion, which adaptively amplifies expressiveness by incorporating semantic-aware key frames into the rhythm-related base motion.
@@
  • We propose SemTalk, a novel framework for holistic co-speech motion generation that separately models rhythm-related base motion and semantic-aware sparse motion, adaptively integrating them via a learned semantic gate.
  • We propose a hierarchical coarse2fine cross-attention module to refine base motion and a local-global rhythmic consistency learning to integrate latent face and hand features with rhythm-related priors, ensuring coherence and rhythmic consistency. We then propose semantic emphasis learning to generate semantic gestures at certain frames, enhancing semantic-aware sparse motion.
@@ 2 Related Work
Co-speech Gesture Generation. Co-speech gesture generation aims to produce gestures aligned with speech. Early rule-based methods [21, 22, 7, 19, 41] lacked variability, while deterministic models [49, 46, 5, 7, 36, 29] mapped speech directly to gestures. Probabilistic models, including GANs [17, 40, 1] and diffusion models [53, 2, 47, 10], introduced variability. Some methods incorporated semantic cues, such as HA2G [32] and SEEG [28], which used hierarchical networks and alignment techniques. SynTalker [8] employs prompt-based control but treats inputs as signal strengths rather than fully interpreting semantics. LivelySpeaker [52] combines rhythmic features and semantic cues using CLIP [39] but struggles to integrate gestures with rhythm and capture semantics consistently, moreover, it only provides global control, limiting fine-grained refinement. DisCo [29] disentangles content and rhythm but lacks explicit modeling of sparse semantic gestures. SemTalk addresses this by separately modeling rhythm-related base motion and semantic-aware sparse motion, integrating them adaptively through a learned semantic score.
Holistic Co-speech Motion Generation. Generating synchronized, expressive full-body motion from speech remains challenging, especially in coordinating the face, hands, and torso [34, 48, 31, 9, 37]. Early methods introduced generative models to improve synchronization, but issues persisted. TalkSHOW [48] improved with VQ-VAE [42] cross-conditioning but handled facial expressions separately, causing fragmented outputs. DiffSHEG [9] and EMAGE [31] used separate encoders for expressions and gestures, but their unidirectional flow limited coherence. ProbTalk [33] leverages PQ-VAE [43] for improved body-facial synchronization but mainly relies on rhythmic cues, risking the loss of nuanced semantic gestures. Inspired by TM2D [15], which decomposes dance motion into music-related components, we separately model co-speech motion into rhythm-related and semantic-aware motion.
@@ 3.1 Preliminary on RVQ-VAE
Following [51, 4, 16], our approach uses a residual vector-quantized autoencoder (RVQ-VAE) to progressively capture complex body movements in a few players. To retain unique motion characteristics across body regions, we segment the body into four parts—face, upper body, hands, and lower body—each with a dedicated RVQ-VAE, following [3, 31]. This segmentation preserves each part’s dynamics and prevents feature entanglement.
@@ 3.2 Overview
As shown in Figure 2, our SemTalk pipeline includes two main components: the Base Motion Blocks $f_{r}(\cdot)$ and the Sparse Motion Blocks $f_{b}(\cdot)$. Given rhythmic features $\gamma_{b}$, $\gamma_{h}$, a seed pose $\tilde{m}$, and a speaker ID $id$, the Base Motion Blocks generate rhythm-aligned codes $q^{b}$, forming the rhythmic foundation of the base motion:
$f_{r}:(\gamma_{b},\gamma_{h},\tilde{m},id;\theta_{f_{r}})\rightarrow q^{b},$ (1)
@@
where $\theta_{f_{s}}$ represents the Sparse Motion Block parameters. Finally, the semantic emphasis mechanism $\mathcal{E}$ combines $q^{b}$ and $q^{s}$, guided by $\psi$, to form the final motion codes $q^{m}$:
$q^{m}=\mathcal{E}(q^{b},q^{s};\psi).$ (3)
@@
Figure 3: Architecture of SemTalk. SemTalk generates holistic co-speech motion in three stages. (a) Base Motion Generation uses rhythmic consistency learning to produce rhythm-aligned codes $q^{b}$, conditioned on rhythmic features $\gamma_{b}$, $\gamma_{h}$. (b) Sparse Motion Generation employs semantic emphasis learning to generate semantic codes $q^{s}$, activated by semantic score $\psi$. (c) Adaptively Fusion automatically combines $q^{b}$ and $q^{s}$ based on $\psi$ to produce mixed codes $q^{m}$ at frame level for rhythmically aligned and contextually rich motions.
@@ 3.3 Generating Rhythm-related Base Motion
The Base Motion Generation (Fig. 3 a) in SemTalk establishes a rhythmically aligned foundation by leveraging both rhythmic and speaker-specific features, enhancing the naturalness and personalization of generated motion.
Rhythmic Speech Encoding. To synchronize motion with speech, SemTalk incorporates rhythmic features: beats $\gamma_{b}$ and HuBERT features $\gamma_{h}$. $\gamma_{b}$, derived from amplitude, short-time energy [11], and onset detection, mark key rhythmic points for aligning gestures with speech. Meanwhile, $\gamma_{h}$, extracted by the HuBERT encoder [18], captures high-level audio traits. In addition to rhythmic features $\gamma$, SemTalk uses a seed pose $\tilde{m}$ and speaker identity $id$ to generate a personalized, rhythm-aligned latent pose $p$. Then MLP-based Face Enhancement and Body Part-Aware modules utilize $\gamma$, $p$ and $id$ to obtain latent face $f_{e}$, hands $f_{h}$, upper body $f_{u}$ and lower body $f_{l}$.
Coarse2Fine Cross-Attention Module. To facilitate the learning of base motion, we first proposed a transformer-based hierarchical Coarse2Fine Cross-Attn Module utilize $f_{e}$, $f_{h}$, $f_{u}$ and $f_{l}$ to obtain latent base motion $f_{b}$. The refinement begins with $\gamma$ for $f_{e}$, which guides the rhythmic representation for $f_{h}$, followed by conditioning $f_{u}$ and finally influencing $f_{l}$. Since mouth movements closely correspond to speech syllables with minimal delay, we use the face to guide hand motions, inspired by DiffSHEG [9]. As the upper and lower body movements are less directly driven by speech and instead reflect the natural swinging of the hands and torso, we adopt cascading guidance: hands influence the upper body, which in turn drives the lower body. This structured approach, moving from the face to the hands, upper body, and lower body, ensures smooth and coherent motion propagation across the entire body.
Rhythmic Consistency Learning. Inspired by CoG’s use of InfoNCE loss [45] to synchronize facial expressions with audio cues, our approach adopts a similar philosophy of aligning motion and speech rhythm. It can be defined as:
@@
Unlike CoG, our approach fundamentally differs by incorporating separate local and global rhythmic consistency losses, which are applied to both latent face $f_{e}$ and latent hands $f_{h}$, ensuring a more cohesive and synchronized representation across the entire motion sequence. This rhythmic consistency loss ensures that the motions are not only synchronized at the frame level but also maintain a consistent rhythmic flow across the entire sequence.
The local frame-level consistency loss $\mathcal{L}_{\text{Rhy}}^{(L)}$ aligns the motion features of each frame with the corresponding rhythmic cues $\gamma_{h}$. By leveraging HuBERT features $\gamma_{h}$ instead of basic beat features $\gamma_{b}$, which only capture rhythmic pauses, we incorporate rich, high-level audio representations that enhance the model’s ability to capture rhythm-related motion patterns and maintain temporal coherence.
The global sentence-level consistency loss $\mathcal{L}_{\text{Rhy}}^{(G)}$ is designed to ensure rhythmic coherence at a global level. Unlike local loss, $\mathcal{L}_{\text{Rhy}}^{(G)}$ reinforces rhythm consistency throughout the sequence, ensuring that the generated motion maintains smooth and rhythm-aligned throughout its duration.
By jointly minimizing $\mathcal{L}_{\text{Rhy}}^{(L)}$ and $\mathcal{L}_{\text{Rhy}}^{(G)}$, rhythmic consistency learning enables SemTalk to produce base motions that are rhythmically aligned and temporally cohesive, forming a solid rhythm-related base motion foundation.